Air enters a diffuser at a constant 109 Kpa , 60 celcius and 292 meters per second. The air exits at 305 Kpa , 10 Celcius and 10 meters per second. The area of the inlet is 104 cm^2. You can assume the air is an ideal gas. CP and CV values are assumed constant. Steady flow. There is no change in Potential Energy.

Ru=8.314 kj/kmol*k

Mair = 28.97 kg/kmol

cp=1.005 kj/kg*k

cv=0.718 kj/kg*k

Determine the average heat transfer time rate in kW.
Answer #1

in SA VIRS= B - 10g 2 O P2 = 305k Pa T2 = 10°c V2 =10m/s. P = 109 kPa TI = 60°c Vi = 292 m/s A = 104 an ² - loy m² (100) 2 1a